OpenStudy (anonymous):

simplify cosX/tan^9(X) - cosx/sin^2(x) please show method for solving

OpenStudy (mertsj):

\[\frac{cosx}{\tan ^{9}x}- \frac{cosx}{\sin ^{2}x}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):